OpenStudy (anonymous):

Find the angle between the vectors u and v if u = - 4j + 3j and v = 3i + 2j. Round answer to two decimal places.

OpenStudy (turingtest):

the angle between two vectors can be found by taking their dot product

OpenStudy (turingtest):

\[\vec u\cdot\vec v=\|\vec u||\vec v\|\cos\theta\]solve for \(\theta\)

OpenStudy (anonymous):

How do I get the dot product? Sorry i'm just really confused on this one

OpenStudy (turingtest):

a dot product for a vector in \(\mathbb R^2\)works as follows:\[(a\hat i+b\hat j)\cdot(c\hat i+d\hat j)=ac+bd\]